Please help me solve for x:
a) 5x(^1/2)=20
b) 8x^-3=64
c) 8(x-2)^3=27
d) (x+3)^sqrt2=25^sqrt8

Answer by edjones(7311)

a)
5x^(1/2)=20
x^(1/2)=20/5
x^(1/2)=4
x=16 square each side.
.
b)
8x^-3=64
1/x^3=64/8
1/x^3=8
8x^3=1 multiply each side by x^3
x^3=1/8
x=1/2 take cube root of each side.
.
c)
8(x-2)^3=27
2(x-2)=3
2x-4=3
2x=7
x=7/2=3.5
.
d)
(x+3)^sqrt(2)=25^sqrt(8)
((x+3)^sqrt2)^sqrt(2)=(25^2sqrt(2))^sqrt(2) Make both sides to the sqrt(2) power.
(x+3)^2=25^4
x+3=25^2 take the sqrt of each side.
x=625-3
x=622
.
Ed